OpenStudy (anonymous):

Find the difference. 9/x^2-6x - 7/x^2-36 WILL BECOME FAN AND WILL GIVE MEDAL FOR BEST ANSWER!!! PLEASE HELP

OpenStudy (anonymous):

@BloomLocke367

OpenStudy (anonymous):

@1289florida

OpenStudy (campbell_st):

you need a common denominator so factor both denomunators \[\frac{9}{x(x-6)} - \frac{7}{(x -6)(x + 6)}\] so to get the common denominator you need \[\frac{9}{x(x-6)} \times \frac{(x + 6)}{(x +6)} - \frac{7}{(x -6)(x + 6)} \times \frac{x}{x}\] whoch becomes \[\frac{9(x + 6) - 7 \times x}{x(x-6)(x+6)}\] simplify the numerator for the answer